The sum of kelly's and bob's age is 29. Kelly's age exceeds bb's by 5 years. How old is each person?

Answer by checkley79(3341)

B+K=29
K=B+5
B+(B=5)=29
2B=29-5
2B=24
B=24/2
B=12 AGE OF BOB.
K=12+5=17 AGE OD KELLY.
PROOF:
12+17=29
29=29